OpenStudy (anonymous):

Find the absolute maximum and minimum values of the following function on the given interval. If there are multiple points in a single category list the points in increasing order in x value. f(x)=(3*ln(x))/2x , [1,3]

OpenStudy (anonymous):

this the function: \(\large f(x)=\frac{3lnx}{2x} = \frac{3}{2}\frac{lnx}{x}\) ???

OpenStudy (anonymous):

in any case, since you're looking for the absolute extrema, you'll need to evaluate at the endpoints of your interval in addition to the x's where f ' =0

OpenStudy (anonymous):

then compare...